OpenStudy (anonymous):

Look at the table of values below. a b 2 8 4 14 6 20 Which equation is represented by the table? B. b = 3a + 2 C. b = 4a - 4 D. b = 4a - 2

OpenStudy (lukecrayonz):

Do you know how to solve these?

OpenStudy (anonymous):

nope @Lukecrayonz

OpenStudy (anonymous):

just need the answr @Lukecrayonz

OpenStudy (lukecrayonz):

Haha can't do that for you now that I know you don't know how to do them! So look (You're missing A) B. b=3a+2 and the chart shows a b 2 8 So put those into the equation for everytime you see a, put 2 instead, and everytime you see b, put 8. 8=3*2+2 8=6+2 8=8 True Now that doesn't mean it's the answer. You have to check every set of numbers.

OpenStudy (anonymous):

got it! C is the answer!!!(" @Lukecrayonz